OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 2x = -1

OpenStudy (anonymous):

reallly stuckkk

OpenStudy (anonymous):

I'm hearing x2 + 26x + ? = 0 where you want to know what the ? would be for a perfect square binomial. I would be remiss if I did not mention that this is not the best method for solving this equation. First, the best method: x2 + 26x = 0 First, just factor the x from the left side: x(x + 26) = 0 Then x = 0 or -26. Problem solved, two answers. If you really want to complete the square, you can still solve it, but it takes a bit longer. Take half of the middle number (26) which is 13 and then square it. 132=169. That is your question mark. Don't forget to add it to the 0 to keep the equation balanced: x2 + 26x + 169 = 0 +169 To finish solving: (x + 13)2 = 169 x + 13 = +/-13 x = -13-13 or x = 13-13 x = -26 or x = 0. Same answer as above. A heckuvalot more complicated. Read more: http://wiki.answers.com/Q/What_number_is_needed_to_complete_the_square_x2_plus_26x_equals_0#ixzz20PvT6Iwh